OpenStudy (rainbow_dash):

log5 (x + 2) - log5 9 = log5 81

OpenStudy (lgbasallote):

shouldnt there be a mathematical operationsomewhere?

OpenStudy (rainbow_dash):

whoops.

OpenStudy (lgbasallote):

\[\LARGE \log_a b - \log_a c \implies \log_ a (\frac{b}{c})\]

OpenStudy (cherylim23):

log5 (x+2) = log5 81 + log5 9 log5 (x+2) = log5 (81*9) x+2 = 729 x=727 hope this is right :)